9.71 An experimental medication is administered to 16 people who suffer from migraines. After an
hour, 10 say they feel better. Is the medication effective (i.e., is the percent who feel better greater
than 50 percent)? Use α = .10, explain fully, and show all steps.

Answer by stanbon(75887)

An experimental medication is administered to 16 people who suffer from migraines. After an hour, 10 say they feel better.
Is the medication effective (i.e., is the percent who feel better greater
than 50 percent)?
Use α = .10, explain fully, and show all steps.
------------------
Ho: p = 0.5
Ha: p > 0.5
---
Critical value for right-tail test with alpha = 10%: z = invNorm(0.90) = 1.2816
test statistic: z(10/16) = (0.625-0.5)/sqrt(0.5*0.5/16) = 1
p-value: P(z>1) = 0.1587
----
Conclusion: Since the p-value is greater than 10%, fail to reject Ho.
At the 10% level of significance the medication is ineffective; the
proportion who feel better is still 50% after medication.
==========================
Cheers,
Stan H.